OpenStudy (anonymous):

What is the simplified form of the expression? (y^–5)^–10y^10 a. y^–60 b. y^60 c. y^–150 d. y^150

OpenStudy (anonymous):

@Zale101

jimthompson5910 (jim_thompson5910):

first focus on simplifying (y^–5)^–10

jimthompson5910 (jim_thompson5910):

how do you do this?

jimthompson5910 (jim_thompson5910):

here's a similar example (x^2)^7 = x^(2*7) = x^(14)

OpenStudy (anonymous):

i dont know either of them..

jimthompson5910 (jim_thompson5910):

notice how I simplified (x^2)^7

jimthompson5910 (jim_thompson5910):

I just multiplied the exponents to get x^(14)

OpenStudy (anonymous):

oh im ctching on now. y^50

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so (y^–5)^–10y^10 turns into y^50*y^10

jimthompson5910 (jim_thompson5910):

how would you simplify y^50*y^10

OpenStudy (anonymous):

y^ 50+10 = y^60

jimthompson5910 (jim_thompson5910):

nailed it